OpenStudy (anonymous):

The original purchase price of a car is $16,000. Each year, its value depreciates by 10%. Three years after its purchase, what is the value of the car? A. $16 B. $34,587.23 C. $11,664 D. $12,960 **not quite sure how to solve this... please explain? :) thx!

jimthompson5910 (jim_thompson5910):

Final Value = Original Value*(100% - 10%)^(Number of years) FV = 16000*(1 - 0.10)^(3) FV = 16000*(0.90)^(3) I'll let you finish

OpenStudy (anonymous):

kk so i get 11664 ? so my answer is C. $11,664 ??

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

awesome!! Thank youu :)